# CLOVA: A Closed-LOop Visual Assistant with Tool Usage and Update

### v1

@@ 3.4 Learning
Method GQA NLVRv2 Image Editing Tagging
End-to-end Otter [24] $48.2$ $48.2$
@@
VISPROG [11] $49.8$ $60.8$ $40.2$ $39.3$
CLOVA (Ours) $\mathbf{54.6}$ $\mathbf{65.6}$ $\mathbf{65.4}$ $\mathbf{50.2}$
@@
Table 3: Comparisons in the four tasks. We report accuracies on GQA, NLVRv2, and image editing tasks, and F1 score on the tagging task.
Prompt training and validation. Given the collected data, we invoke training-validation prompt tuning to update tools. Note that, instead of training a single soft prompt for all the collected data, we choose to learn a soft prompt for each training instance collected. Each learned prompt will then be validated by running the tool with it on validation data (held out from collected data except for the VQA tool, where VQA will be validated on the original visual question it failed on) and seeing if the tool can produce the response we expect in an update (e.g. correctly localizing a horse for the LOC tool). As a result, we can discard prompts that do not lead to the expected tool update possibly due to the faulty training instances they were trained on, alleviating the issue of possibly noisy collected data. Finally, we build a prompt pool $\mathcal{P}$ for each tool. Taking the LOC tool as an example: after training and validation, CLOVA stores the visual concept (e.g., “horse” in the task of localizing horses), the learned prompts and the visual features of all collected instances in $\mathcal{P}$, formulated as $\mathcal{P}=\left\{v_{j}:\big{[}[f_{j1},\cdots,f_{jn}],[p_{j1},\cdots,p_{jn}]\big{]}\right\}_{j=1}^{m}$ where $v_{j}$ is the name of the $j$-the concept, $m$ concepts are stored totally. For concept $v_{j}$, $f_{ji}$ and $p_{ji}$ are the feature and learned prompt for the $i$-th instance, respectively, and $n$ instances are learned for $v_{j}$.
Prompt ensemble. During inference, we use prompt ensemble to retrieve and utilize prompts from the aforementioned prompt pool $\mathcal{P}$. Specifically, given an instruction and the generated program, we first identify the visual concept of each tool being used. For example, given an image editing task, “Replace the dog with a cat” where SEG, SELECT, and REPLACE models will be used, “dog” is the visual concept for SEG and SELECT, and “cat” is the visual concept for REPLACE. Then, in each step of tool-use with an input image $x$, if the visual concept of the tool-use is not in $\mathcal{P}$ of the corresponding tool, the soft prompt $p^{\prime}$ will be set as a zero-vector, i.e. the tool has not been updated with this concept so we just use the original tool; if it can be found in $\mathcal{P}$, i.e. the tool was updated with $v_{j}$ before, we aggregate the soft prompt corresponding to this concept based on the similarity between the visual features stored with these prompts and feature extracted from the current input image $x$: $p^{\prime}$ is computed by $p^{\prime}=\frac{\sum_{i=1}^{n}w_{i}\cdot p_{ji}}{\sum_{i=1}^{n}w_{i}}$, where we compute the cosine similarity between feature $f_{x}$ of $x$ and features $f_{ji}$ in $\mathcal{P}$ as the weight $w_{i}$.
3.4.1 LLM learning with demonstrations
Besides the tools, CLOVA can also update its LLMs through in-context learning. As we mentioned above, CLOVA utilizes a demonstration pool $\mathcal{D}$ to provide relevant examples for the LLMs. After working on a new problem, the plan, program, and reflection will be stored into $\mathcal{D}$ based on whether this problem is correctly solved (then it will be marked as a correct example) or the opposite (as an incorrect example). We also have a validation process that uses the original instruction as validation data to evaluate stored in-context examples. As the size of $\mathcal{D}$ grows, LLMs use more examples and therefore strengthen reasoning skills.
The framework of CLOVA is summarized in Algorithm 1.
@@ 3.4.1 LLM learning with demonstrations
0:  LLMs, visual models, instruction data $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{t}\}$, demonstration pool $\mathcal{D}$, prompt pool $\mathcal{P}=\emptyset$.
@@
8:        Perform training-validation prompt tuning, store in-context examples into $\mathcal{D}$ and soft prompts in $\mathcal{P}$.
9:        if Updated models solve $T_{i}$ incorrectly then
@@
11:           Perform training-validation prompt tuning, store in-context examples into $\mathcal{D}$ and soft prompts in $\mathcal{P}$.
@@
4 Experiments
4.1 Setting
To comprehensively evaluate the learning ability of CLOVA, we set a separate training stage before deployment, which iteratively learns knowledge via inference, reflection, and updating phases. In the test stage (i.e., after development), we do not update LLMs and visual models, and evaluate the performance only via the inference phase.
Similar to VISPROG, we manually collect data for the language-guided image editing and factual knowledge tagging tasks. To better evaluate the learning capability, we collect fine-grained visual concepts that visual models may not have learned, such as “Replace the lion in the image with pine grosbeak”, where pine grosbeak is a fine-grained bird of Passeriformes. In the image editing task, we collect $129$ images with $193$ instructions, where $27$ images with $78$ instructions are used for training, and the rest are test data. We manually check whether edited images are semantically correct. The factual knowledge tagging task needs the model to identify persons or objects with bounding boxes in images. We collect $86$ images with $177$ instructions for this task, where $10$ images with $88$ instructions are used for training and the rest are used as the test data. We report the F1 score for this task.
In plan and program generation, prompts contain $4$ correct examples and $4$ incorrect examples. The demonstration pool $\mathcal{D}$ is initialized having about $15$ in-context examples.
Figure 6: Case study of CLOVA on four example tasks.
4.2 Main Results
We compare CLOVA with tool-usage methods: VISPROG [11], GPT4TOOLs [72], Visual ChatGPT [67], InternGPT [30], HuggingGPT [53], and ViperGPT [58]. We use their official codes, where all methods use the GPT-3.5 model. In addition, we also compare CLOVA with two advanced end-to-end models: Otter [24] and MMICL [76], which do well in GQA and NLVRv2 datasets. Results on the four tasks are shown in Tab. 3. We observe that CLOVA achieves the best performance among tool-usage methods. CLOVA has at least $4.8\%$ improvements on the GQA dataset and $4.9\%$ improvements on the NLVRv2 dataset. The reason is that CLOVA learns how to generate programs for the two tasks and update the VQA and LOC models for better image perception. CLOVA performs competitively and even outperforms Otter and MMICL.
On image editing and knowledge tagging tasks, we do not compare our method with InterGPT, HuggingGPT, and ViperGPT, since they either need to manually input object masks or cannot accurately locate fine-grained objects. In addition, most methods cannot finish tagging tasks. Thus, we compare our method with VISPROG and Visual ChatGPT. As we collect find-grained data, it is challenging for off-the-shelf classification, segmentation, and image generation models. Since GPT4TOOLs and Visual ChatGPT cannot use OpenCV functions and do not have the learning capability, they get bad performance on the image editing task. VISPROG can use OpenCV functions, but it cannot learn new knowledge. Its main fault is the inability to recognize or generate fine-grained concepts. Compared with them, the learning ability of CLOVA brings more than $20\%$ improvements in image editing tasks, and $10\%$ improvements in knowledge tagging tasks.
4.3 Case Study
In Fig. 6, we visualize four cases to illustrate the reflection and learning ability in CLOVA. It is capable of identifying models that need to be updated, no matter LLMs or visual models. Hard prompts guide LLMs to generate correct programs for similar instructions. Visual models could learn new concepts via the proposed data collection manners and soft prompt turning scheme. In Fig. 7, we visualize an example of global and local reflection. When the instruction is complex, the global reflection does not accurately identify which step has the error. Using global reflection as in-context examples still cannot generate correct programs. In contrast, local reflection successfully identifies the error step and using local reflection generates correct programs.
Method GQA NLVRv2
Reflection w/o local reflection $52.0$ $65.2$
w/o global reflection $53.6$ $64.2$
w/o intermediate results $48.8$ $61.2$
w/o plan $50.0$ $62.6$
Ours $\mathbf{54.6}$ $\mathbf{65.6}$
Prompt Tuning for LLMs w/o incorrect cases $46.1$ $61.4$
w/o correct cases $48.2$ $63.2$
w/o validation $44.2$ $61.0$
Ours $\mathbf{54.6}$ $\mathbf{65.6}$
Prompt Tuning for visual models w/o validation $42.8$ $62.8$
Ours $\mathbf{54.6}$ $\mathbf{65.6}$
Table 4: Ablation on the GQA and NLVRv2 dataset.
4.4 Ablation
We conduct ablations to evaluate the reflection and learning phases, using the GQA and NLVRv2 datasets. For reflection, we evaluate only using global reflection, only using local reflection, not using multimodal intermediate results, and not generating plans. We separately evaluate hard and soft prompts for learning. We evaluate only storing correct examples and only storing incorrect examples for hard prompt tuning. We also evaluate removing the validation process in hard and soft prompt tuning processes. Results are shown in Tab. 4. We find that these components are necessary for CLOVA to achieve better performance.
We evaluate our method using different LLMs: Llama2-7B, GPT-3.5-turbo, and GPT-4. The versions GPT-3.5-turbo and GPT-4 are in October 2023. Results on GQA and NLVRv2 are shown in Tab. 5. We find that on both two datasets with three LLMs, updating LLMs and visual models indeed leads to improvements. In addition, we observe that our method achieves higher improvements on open-source LLMs (i.e., Llama2-7B), $21\%$ on the GQA dataset and $13.8\%$ on the NLVRv2 dataset, bringing significance of studying the learning capability of visual assistants.
Dataset Method LLama2-7B GPT-3.5-turbo GPT-4
GQA Baseline $39.2$ $46.4$ $52.6$
+ Update LLMs $56.8$ $51.6$ $56.6$
+ Update visual models $60.2$ $54.6$ $60.4$
NLVRv2 Baseline $50.0$ $60.2$ $64.8$
+ Update LLMs $59.2$ $63.6$ $68.8$
+ Update visual models $63.8$ $65.6$ $69.2$
Table 5: Different LLMs on the GQA and NLVRv2 datasets.
5 Conclusion
In this paper, we have presented CLOVA, a general visual assistant that can adapt to new environments via inference, reflection, and learning in a closed-loop framework. In the inference phase, using both correct and incorrect examples for prompts benefits to generate better plans and programs. Our reflection scheme is capable of identifying tools that need to be updated. Through three flexible data collection manners and the validation-learning prompt tuning scheme in the learning phase, CLOVA can efficiently improve its tools while avoiding catastrophic forgetting. Experimental results on four tasks show the effectiveness of CLOVA as a general visual assistant with learning abilities. In the current method, we assume there is no selection or loop structure in programs, and assume there is at most one model that needs updates in a task. The two assumptions cannot always hold in real world. We are going to break the assumptions, making the closed-loop framework more practical.
Supplementary Material
@@ 6 Comparisons with related methods
In Tab. 6, we present a more detailed comparison table with more related methods. In the table, ‘Global’ means the global reflection, ‘Local’ means the local reflection, ‘Instruction’ means instruction-following tuning, ‘RL’ means reinforcement learning, and ‘Prompt’ means using prompt examples as in-context learning. We observe that many tool-usage methods do not have a reflection ability, and few methods use global reflection to improve the plan or programs generated by LLMs. Different from them, CLOVA uses both global reflection and local reflection to identify models that need to be updated, capable of handling complex instructions. Moreover, as we all know, our method is the first work to update visual tools, through which the visual assistant can better adapt to new environments.
@@
Table 6: Comparisons with representative tool-usage methods, where VMs means visual models.

### v2

@@ 3.4.1 Updating tools with prompt tuning
Prompt tuning and validation. Given the collected data, we invoke training-validation prompt tuning to update tools. Note that, instead of learning a single prompt for all collected data, we choose to learn a prompt for each training instance collected. Each learned prompt will then be validated by running the tool with it on validation data (held out from collected data except for the VQA tool, where VQA will be validated on the original visual question it failed on) and seeing if the tool can produce desirable responses (e.g. correctly localizing a horse for the LOC tool). As a result, we discard prompts that do not lead to the desirable responses, possibly due to the faulty training instances they were trained on, alleviating the issue of noisy collected data. Finally, we build a prompt pool $\mathcal{P}$ for each tool. Take the LOC tool as an example. After training and validation, CLOVA stores the visual concept (e.g., “horse” in the task of localizing horses) with its learned prompts and visual features of all collected instances in $\mathcal{P}$, formulated as $\mathcal{P}=\left\{v_{j}:\big{[}[f_{j1},\cdots,f_{jn}],[p_{j1},\cdots,p_{jn}]\big{]}\right\}_{j=1}^{m}$, where $v_{j}$ is the name of the $j$-the concept (e.g., “horse”), $m$ concepts are stored totally. For the concept $v_{j}$, $f_{ji}$ and $p_{ji}$ are the feature and learned prompt using the $i$-th instance, respectively, and $n$ instances are learned for $v_{j}$.
In summary, a tool is formulated as $T_{\theta,\mathcal{P}}$, where $\theta$ is the parameter of neural networks. The forward process of a sample $x$ is $T_{\theta,\mathcal{P}}(x)=\theta([x,\mathcal{P}])$, where we concatenate $x$ with a retrieved prompt from $\mathcal{P}$ as the input for the tool. We update the tool $T_{\theta,\mathcal{P}}$ by tuning $\mathcal{P}$ while $\theta$ being fixed,
$\min_{\mathcal{P}}\mathbb{E}_{(x,y)}\mathcal{L}\big{(}T_{\theta,\mathcal{P}}(x),y\big{)},$
where $(x,y)$ is collected data, $\mathcal{L}$ is the loss function of $T_{\theta,\mathcal{P}}$.
Prompt ensemble. During inference, we use prompt ensemble to retrieve and utilize prompts from the learned prompt pool $\mathcal{P}$. Specifically, given a generated program, we first identify the visual concept for each involved tool in the program. For example, given an image editing task, “Replace the dog with a cat” where the SEG, SELECT, and REPLACE tools will be used, “dog” is the visual concept for the SEG and SELECT tools, and “cat” is the visual concept for the REPLACE tool. Then, in each step of tool usage with an input image $x$, if the visual concept is not in $\mathcal{P}$ of the corresponding tool, the prompt $p^{\prime}$ for $x$ will be set as a zero-vector, i.e. the concept has not been learned for the tool so we just use the original tool (using zero-vector as a prompt); if the visual concept can be found in $\mathcal{P}$, i.e. the tool was updated with the visual concept before, we aggregate the prompts corresponding to this concept based on the similarity between features stored with these prompts and the feature extracted from the current input $x$: $p^{\prime}$ is computed by $p^{\prime}=\frac{\sum_{i=1}^{n}w_{i}\cdot p_{ji}}{\sum_{i=1}^{n}w_{i}}$, where we compute the cosine similarity between feature $f_{x}$ of $x$ and features $f_{ji}$ in $\mathcal{P}$ as the weight $w_{i}$.
3.4.2 Updating LLMs with demonstrations
Besides the visual tools, CLOVA can also update its LLMs. As we mentioned above, CLOVA utilizes a demonstration pool $\mathcal{D}$ to provide relevant examples for the LLMs. After working on new data, the plan, program, and reflection will be stored in $\mathcal{D}$ as correct or incorrect examples, based on whether the data is correctly proceeded. We also have a validation process that uses the original instruction as validation data to evaluate stored in-context examples. As the size of $\mathcal{D}$ grows, LLMs use more examples and therefore strengthen reasoning skills.
4 Experiments
4.1 Setting
To comprehensively evaluate the learning capability of CLOVA, we set a separate training stage before deployment, which iteratively learns new knowledge via inference, reflection, and updating phases. In the test stage (i.e., after development), we do not update LLMs and visual tools, and evaluate the performance only via the inference phase.
Similar to VISPROG, we manually collect data for the language-guided image editing and factual knowledge tagging tasks. To better evaluate the learning capability, we collect fine-grained visual concepts that visual tools may not have learned, such as “Replace the lion in the image with pine grosbeak”, where pine grosbeak is a fine-grained bird of Passeriformes. In the image editing task, we collect $129$ images with $193$ instructions, where $27$ images with $78$ instructions are used for training, and the rest are test data. We manually check whether edited images are semantically correct. The factual knowledge tagging task needs to identify persons or objects with bounding boxes in images. We collect $86$ images with $177$ instructions for this task, where $10$ images with $88$ instructions are used for training and the rest are used as the test data. We report the F1 score for this task.
In plan and program generation, prompts contain $4$ correct examples and $4$ incorrect examples. The demonstration pool $\mathcal{D}$ is initialized having about $20$ in-context examples.
Figure 6: Case study of CLOVA on four example tasks.
4.2 Main Results
Method GQA NLVRv2 Editing Tagging
E2E Otter [26] $48.2$ $48.2$
@@ 4.2 Main Results
VISPROG [10] $49.8$ $60.8$ $40.2$ $0.393$
CLOVA (Ours) $\mathbf{54.6}$ $\mathbf{65.6}$ $\mathbf{65.4}$ $\mathbf{0.502}$
@@
Table 3: Comparisons in the four tasks. We report accuracies on GQA, NLVRv2, and image editing tasks, and F1 score on the knowledge tagging task. ‘E2E’ means end-to-end methods.
We compare CLOVA with tool-usage methods: VISPROG [10], GPT4TOOLs [78], Visual ChatGPT [73], InternGPT [39], HuggingGPT [61], and ViperGPT [66]. We use their official codes, where all methods use the GPT-3.5 model. In addition, we also compare CLOVA with two advanced end-to-end models: Otter [26] and MMICL [81], which do well in GQA and NLVRv2 datasets. Results on the four tasks are shown in Tab. 3. We observe that CLOVA achieves the best performance among tool-usage methods. CLOVA has at least $4.8\%$ improvements on the GQA dataset and $4.9\%$ improvements on the NLVRv2 dataset. The reason is that CLOVA learns how to generate programs for the two tasks and update the VQA and LOC tools for better image perception. CLOVA performs competitively and even outperforms Otter and MMICL.
On image editing and knowledge tagging tasks, we do not compare our method with InterGPT, HuggingGPT, and ViperGPT, since they either need object masks or cannot accurately locate objects. In addition, most methods cannot finish the tagging task. Thus, we compare our method with VISPROG and Visual ChatGPT. As we collect find-grained data, it is challenging for off-the-shelf classification, segmentation, and image generation tools. Since GPT4TOOLs and Visual ChatGPT cannot use OpenCV functions and do not have the learning capability, they get bad performance on the image editing task. VISPROG can use OpenCV functions, but it cannot learn new knowledge. Its main fault is the inability to recognize or generate fine-grained concepts. Compared with them, the learning capability of CLOVA brings more than $20\%$ and $10\%$ improvements to image editing and knowledge tagging tasks, respectively.
4.3 Qualitative Results
In Fig. 6, we visualize four cases to illustrate the reflection and learning capability in CLOVA. It identifies tools that need to be updated, no matter LLMs or visual tools. Prompt engineering guides LLMs to generate correct programs for similar instructions. Visual tools learn new concepts via our data collection and prompt turning schemes. In Fig. 7, we visualize an example of global and local reflection. When the instruction is complex, the global reflection does not accurately identify which step has the error. Using global reflection as in-context examples still cannot generate correct programs. In contrast, local reflection successfully identifies the error step, and using local reflection generates correct programs, showing the effectiveness of local reflection.
Method GQA NLVRv2
Reflection w/o local reflection $52.0$ $65.2$
w/o global reflection $53.6$ $64.2$
w/o intermediate results $48.8$ $61.2$
w/o plan $50.0$ $62.6$
Ours $\mathbf{54.6}$ $\mathbf{65.6}$
Prompt Engineering for LLMs w/o incorrect cases $46.1$ $61.4$
w/o correct cases $48.2$ $63.2$
w/o validation $44.2$ $61.0$
Ours $\mathbf{54.6}$ $\mathbf{65.6}$
Prompt Tuning for visual tools w/o validation $42.8$ $62.8$
Ours $\mathbf{54.6}$ $\mathbf{65.6}$
Table 4: Ablation on the GQA and NLVRv2 dataset.
4.4 Ablation Studies
We conduct ablation studies on the reflection and learning phases, using the GQA and NLVRv2 datasets. For reflection, we evaluate only using global reflection, only using local reflection, not using multimodal intermediate results, and not generating plans. We separately evaluate learning schemes for LLMs and visual tools. We evaluate only storing correct examples or incorrect examples for updating LLMs. We also evaluate removing the validation process in prompt engineering and prompt tuning processes. Results are shown in Tab. 4. We find that these components are necessary for CLOVA to achieve better performance.
We evaluate CLOVA using different LLMs: LLaMA2-7B, GPT-3.5-turbo, and GPT-4. Results on GQA and NLVRv2 are shown in Tab. 5. We find that CLOVA leads to improvements in both strong LLMs (GPT-4) and weaker LLMs (GPT-3.5 and LLaMA2-7). We observe that CLOVA even achieves higher improvements on open-source LLMs (i.e., LLaMA2-7B), $21\%$ on the GQA dataset and $13.8\%$ on the NLVRv2 dataset, bringing the significance of studying the learning capability of visual assistants. We further conduct experiments to evaluate CLOVA on two open-source LLMs: LLaMA2-7B and Mistral-7B. Results are shown in Tab. 6. We observe that CLOVA achieves significant improvements again.
Dataset Method LLaMA2-7B GPT-3.5 GPT-4
GQA Baseline $39.2$ $46.4$ $52.6$
+ Update LLMs $56.8$ $51.6$ $56.6$
+ Update visual tools $60.2$ $54.6$ $60.4$
NLVRv2 Baseline $50.0$ $60.2$ $64.8$
+ Update LLMs $59.2$ $63.6$ $68.8$
+ Update visual tools $63.8$ $65.6$ $69.2$
Table 5: Different LLMs on the GQA and NLVRv2 datasets.
Method GQA NLVRv2 Editing Tagging
LLama2-7B $39.2$ $50.0$ $31.2$ $0.308$
LLama2-7B + Ours $\mathbf{60.2}$ $\mathbf{63.8}$ $\mathbf{47.6}$ $\mathbf{0.357}$
Mistral-7B $20.4$ $34.6$ $29.0$ $0.205$
Mistral-7B + Ours $\mathbf{31.4}$ $\mathbf{42.2}$ $\mathbf{46.5}$ $\mathbf{0.303}$
Table 6: Results on two open-source LLMs.
5 Conclusion and Future Work
In this paper, we have presented CLOVA, a general visual assistant that can adapt to new environments via inference, reflection, and learning in a closed-loop framework. In the inference phase, using both correct and incorrect examples for prompts benefits to generate better plans and programs. Our reflection scheme is capable of identifying tools that need to be updated. Through three data collection manners and the validation-learning prompt tuning scheme in the learning phase, CLOVA can efficiently improve its tools. Experimental results on four tasks and different LLMs show the effectiveness of CLOVA as a general visual assistant with learning abilities.
In the current method, we assume there is no selection or loop structure in programs, and assume there is at most one tool that needs updates in a task. The two assumptions cannot always hold in the real world. We could add selection and loop in-context examples for program generation and iterate the reflection and learning phases to update multiple tools. Besides, we could make some deployment designs to save the response time in our loop, including a foreground process and a background process. The former performs inference and gathers human feedback, while the latter does reflection, updates tools, and periodically synchronizes the weights of tools.
The framework of CLOVA can be easily generalized to new tools. (1) We will try multimodal LLMs (e.g., LLaVA-1.5 [36]) to replace LLMs. In this case, we will evaluate the effectiveness of visual information for plan and program generation, reflection, and answer inference. (2) We can add more up-to-date tools (e.g., BLIP2 [29] as a VQA tool and SAM [25] as an SEG tool), by just designing and programming their forward and prompt tuning processes.
Acknowledgements. We thank the anonymous reviewers for their constructive suggestions. Their insights have greatly improved the quality and clarity of our work. This work was supported in part by the National Science and Technology Major Project (2022ZD0114900).
Supplementary Material
6 Framework of CLOVA
The pseudo-code of CLOVA is summarized in Algorithm 1.
@@ 6 Framework of CLOVA
0:  LLMs, visual tools, instruction data $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{t}\}$, demonstration pool $\mathcal{D}$, prompt pool $\mathcal{P}=\emptyset$.
@@
8:        Perform training-validation prompt tuning, store in-context examples into $\mathcal{D}$ and prompts in $\mathcal{P}$.
9:        if Updated tools solve $T_{i}$ incorrectly then
@@
11:           Perform training-validation prompt tuning, store in-context examples into $\mathcal{D}$ and prompts in $\mathcal{P}$.
@@ 7 Comparisons with related methods
In Tab. 7, we present a more detailed comparison table with more related methods. In the table, ‘Global’ means the global reflection, ‘Local’ means the local reflection, ‘Instruction’ means instruction-following tuning, ‘RL’ means reinforcement learning, and ‘Prompt’ means using prompt examples as in-context learning. We observe that many tool-usage methods do not have a reflection capability, and few methods use global reflection to improve the plans or programs generated by LLMs. Different from them, CLOVA uses both global reflection and local reflection to identify tools that need to be updated, capable of handling complex instructions. Moreover, as we all know, our method is the first work to update visual tools, through which the visual assistant can better adapt to new environments.
@@
Table 7: Comparisons with representative tool-usage methods.